OpenStudy (anonymous):

The cruise control of Megan's truck is set at 65 miles per hour, but her actual speed may vary by as much as 5 miles per hour. Write an inequality that can be used to determine the actual speeds at which Megan could be traveling.

OpenStudy (johnweldon1993):

okay...so it is at 65 right? it cant go below 60...and it cant go above 70 correct? what would that look like?

OpenStudy (anonymous):

I have no idea man

OpenStudy (johnweldon1993):

lol hint.....60 < 65 correct? this is saying....65 is greater than 60 65 < 70 this is saying 70 is greater than 65 right? if you cannot exceed 70 mph...or fall below 60....your equation would look like \[60 \le 65 \le 70\] see why? you cannot go above 70...but you CAN reach 70 you cannot go below 60...but you CAN reach 60 do you agree?

OpenStudy (anonymous):

the answer choices are: A. |s – 65| > 5 B. \[|s – 65| \ge 5 C. |s – 65| < 5 D. |s + 65| \le 5\]

OpenStudy (johnweldon1993):

ahh okay....so....s being speed well you can use the same logic to think about it,..but the answer would be C

OpenStudy (anonymous):

thank you!!! :D

OpenStudy (johnweldon1993):

no problem :)